OpenStudy (anonymous):

Find the absolute maximum for the function f(x)=1/2x+cosx on [-pi/2, pi/2]

OpenStudy (anonymous):

what your want to do here is find first derivation of f(x): f(x) = 1/2x + cos x f'(x) = 1/2 - sin x now, you want to find critical points of your graph, and you do that by setting the first derivate to be equal to zero 0 = 1/2 - sin x -1/2 = - sin x sin x = 1/2 x = sin^-1 (1/2) x = pi/6 we know that x-coordinate of maximum in this case must be greater then 0, because given interval is between [-pi/2, pi/2], and because of 1/2x in the original equation, if we insert a negative number, we can't reach our maximum. with this, we can state that pi/6 is maximum on interval [-pi/2, pi/2]